# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v2

@@ Abstract
This paper studies a class of mean-field control (MFC) problems with singular control under general dynamic state-control-law constraints. We first propose a customized relaxed control formulation to cope with the dynamic mixed constraints and establish the existence of an optimal control using compactification argument in the proper canonical spaces to accommodate the singular control. To characterize the optimal pair of regular and singular controls, we treat the controlled McKean-Vlasov process as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on canonical spaces with constraints on Banach space, allowing us to derive the stochastic maximum principle (SMP) and a constrained BSDE using a novel Lagrange multipliers method. Additionally, we investigate the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control.
Keywords: Mean-field control, singular control, dynamic state-control-law constraint, compactification, stochastic maximum principle, Lagrange multiplier, constrained FBSDE
2020 MSC: 60H30, 49N80, 46N10, 60H10
@@ 1 Introduction
To overcome the high complexity in large stochastic system caused by the mutual influence among agents, mean-field game (MFG) problems (see [22] and [26]) has been proposed and popularized in academic studies, which is to find a Nash equilibrium in a competitive game formulation where each agent’s impact on the aggregate distribution is negligible. MFC problems has also gained considerable attention, which concerns the cooperative population where all agents share the same goal to attain the social optimum. The present paper targets at the MFC problem, which is often interpreted as an optimal control of McKean-Vlasov dynamics by a social planner. Some pioneer studies on this topic, just to name a few, can be found in [2, 7, 10, 1, 5] for the SMP method; [25] and [15] for the limit theory; [27] and [30] for the DPP and HJB equation on the Wasserstein space of probability measures; [4], [31] and [32] for the viscosity solution and comparison principle of the HJB equation. However, it is notable that most existing studies on MFC are limited to regular controls.
@@
The goal of the paper is to contribute to the study of MFC problems by featuring both the singular control and the joint dynamic state-control-law constraints. Motivated by practical risk regulations, liquidity requirements, or feasibility restraints, different types of constraints have been considered in stochastic control problems, but predominantly in single agent’s model. Recently, in mean-field models with only regular controls, there are some emerging research to incorporate state and/or law constraints, see [18] for the level-set approach to tackle the law constraint in MFC problems; [12] for the analysis of mean-field limits under state constraints; [13] for the formulation of controlled FP equation under state constraints; [6] for the optimization formulation of the constrained MFC problem and the derivation of the SMP using the generalized Lagrange multipliers method.
@@
The present paper fills the gap in the literature by advancing some theoretical foundations for constrained MFC problems with singular control. The first contribution of the paper is to address the existence of an optimal pair of regular and singular controls by developing the customized relaxed control formulation and martingale problem to embed the dynamic mixed constraints. That is, we embed the dynamic mixed constraints in item (iv) of the admissible control rules in Definition 3.1. We then present in Lemma 3.5 a martingale problem characterization associated to the admissible control rules and establish in Lemma 3.7 the closedness property of the set of probability measures fulfilling the constraints. By developing technical compactification arguments to accommodate the singular control and combining the previous lemmas using the embedded constraints, we are able to show that the set of approximately optimal control rules is compact (see Lemma 3.8), and hence the set of optimal control rules is compact and convex that yields the existence of an optimal relaxed control (see Theorem 3.9). By standard argument, we also obtain the existence of an optimal strict control for the constrained MFC problem with singular control.
@@
Our second contribution is to further characterize the optimal controls by establishing the SMP for the constrained MFC problem with singular control and derive the adjoint constrained BSDE. However, some conventional techniques such as the first order variation or the spike variation of controls may fail in our setting as the mixed constraints might be violated. Inspired by our recent study [6] for constrained MFC with regular controls, we propose to regard the controlled McKean-Vlasov dynamics with singular control as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on some proper canonical spaces with general constraints on Banach space (see the formulation (4)-(17)). Similar to [6], we are allowed to derive the SMP by using the generalized Fritz-John (FJ) optimality condition and the Lagrange multipliers method. To be more precise, when the forward McKean-Vlasov SDE is taken as an equality constraint, the FJ condition for the control variable gives the SMP, and the Lagrange multipliers in the space of Itô processes induced by the FJ condition on the state variable yields the adjoint processes as the solution to the constrained BSDE, see Theorem 4.2.
As the third contribution of the present paper, we also analyze the non-standard constrained FBSDE in (5)-(27) associated to our constrained MFC problem with singular control. It is worth noting that the Lagrange multiplier $\eta$ in Theorem 4.2 is treated as part of the solution to the constrained BSDE problem. We then verify the constraints qualification and obtain the KKT optimality condition in Lemma 5.1. As a result, we further establish the uniqueness of the solution to the nonstandard constrained FBSDE in (5)-(27), see Theorem 5.4. Furthermore, under some enhanced assumptions, we also prove the stability result of the solution in the sense of (5.6) in Theorem 5.6.
@@
The rest of the paper is organized as follows. Section 2 introduces the constrained MFC problem with singular control and standing assumptions imposed on the model. Section 3 gives the relaxed control formulation and establishes the existence of optimal control to the constrained MFC problem by applying compactification arguments. Section 4 formulates the equivalent optimization problem with constraints in Banach space, for which we develop the SMP and ajoint constrained BSDE using Lagrange multipliers method. Lastly, the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control are discussed in Section 5. The proofs of some auxiliary results are provided in Appendix.
@@ 2.1 MFC with singular control and mixed constraints
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. For $r,l,k\in\mathbb{N}$ and $p>2$, let $W=(W_{t})_{t\in[0,T]}$ be a standard $r$-dimensional $(\mathbb{P},\mathbb{F})$-Brownian motion and $U\subset\mathbb{R}^{l}$ be a closed and convex (regular) set as the policy space. Define the space of regular controls $\mathscr{U}[0,T]$ as the set of $\mathbb{F}$-progressively measurable processes such that $\mathbb{E}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. For the space of singular controls, define $A^{k}[0,T]$ as the set of functions $a:[0,T]\mapsto\mathbb{R}^{k}$ such that $a=(a_{t}^{i})_{t\in[0,T]}$ is càdlàg and non-decreasing with $a_{0}^{i}=0$ for $i=1,\dots,k$. Let $\mathcal{A}^{k}[0,T]$ be the set of $\mathbb{F}$-progressively measurable processes $\zeta=(\zeta_{t})_{t\in[0,T]}$ with paths in $A^{k}[0,T]$ and $\mathbb{E}[|\zeta_{T}|^{p}]<\infty$.
@@
For the above controlled McKean-Vlasov dynamics, we introduce dynamic state-control-law constraints. Let $d\in\mathbb{N}$ be the number of inequality constraints. For $i\in I:=\{1,\dots,d\}$, consider $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $i\in I$ such that the dynamic state-control-law inequality constraints are described by, $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
@@
$\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}J(\alpha,\zeta):=\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}\mathbb{E}\left[\int_{0}^{T}f(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right],$ (3)
where $f:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ and $c:[0,T]\mapsto\mathbb{R}^{k}$ are respectively the running cost function and the cost function associated with singular controls.
@@ 2.2 Standing assumptions
    The mappings $b,\sigma,f$ and $\phi^{i}$, for $i\in I$ are Borel measurable, $b,\sigma$ are continuous and $f$ is lower semi-continuous (l.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$.
@@
    $M(1+|x|^{p}+\sqrt{M_{2}(\mu)})+M_{1}|u|^{p}\leq f(t,x,\mu,u)\leq M(1+|x|^{p}+\sqrt{M_{2}(\mu)}+|u|^{p}).$
@@
    The mappings $b,\sigma,f$ and $\phi^{i}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $\phi^{i}$ is jointly continuous in $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$.
@@
    There exists $M>0$ such that $|\delta(t,x,\mu,u)|\leq M(1+|x|+|u|+\sqrt{M_{2}(\mu)})$ for $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ and $\delta\in\{b,\sigma,\nabla_{u}b(\cdot),\nabla_{u}\sigma(\cdot),\nabla_{x}f(\cdot),\nabla_{u}f(\cdot)\}$.
@@
    The mappings $b(t,x,\mu,u),\sigma(t,x,\mu,u),f(t,x,\mu,u)$ and $\phi^{i}(t,x,\mu,u)$ are continuously Lions-differentiable w.r.t. $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there exists $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\mu}f(t,x,\mu,u)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}\leq M\left(1+|x|+|u|+\sqrt{M_{2}(\mu)}\right),$
@@
    $G:[0,T]\mapsto\mathbb{R}^{n\times k}$ and $c:[0,T]\mapsto[0,\infty)^{k}$ are deterministic continuous functions.
@@
    The constrained function $\phi$ is independent of the control variable and has the linear form $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t)$ for $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Here, $A=(A(t))_{t\in[0,T]}$, $B=(B(t))_{t\in[0,T]}$ and $C=(C(t))_{t\in[0,T]}\in C([0,T];\mathbb{R}^{n})$ and $|A(t)|>0$ and $|A(t)+B(t)|>0$ for $t\in[0,T]$.
@@
    The Hamiltonian $H$ defined by $H(t,x,\mu,u,y,z):=b(t,x,\mu,u)^{\top}y+\mathrm{tr}(\sigma(t,x,\mu,u)^{\top}z)+f(t,x,\mu,u)$ is strictly convex in $u\in U$. There exists a measurable mapping $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ with at most quadratic growth in $(x,\mu,y,z)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ such that $\hat{u}(t,x,\mu,y,z)\in\mathop{\arg\min}_{u\in U}H(t,x,\mu,u,y,z)$ and $\hat{u}$ is Lipschitz continuous, i.e., there is $M>0$ such that, for $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$,
@@
    Define $F(t,x,\mu,u,y,z):=\nabla_{x}H(t,x,\mu,u,y,z)+\int_{\mathbb{R}^{n}}\partial_{\mu}H(t,x^{\prime},\mu,u,y,z)(x)\mu(\mathrm{d}x^{\prime})$,
    $\hat{F}(t,x,\mu,y,z):=F(t,x,\mu,\hat{u}(t,x,\mu,y,z),y,z)$ and $\Phi(t,x,\mu,y,z):=(-\hat{F},\hat{b},\hat{\sigma})(t,x,\mu,y,z)$. The following monotone condition holds that

### v3

# Constrained mean-field control with singular controls: Existence, stochastic maximum principle and constrained FBSDE

@@ Abstract
This paper studies a class of mean-field control (MFC) problems with singular controls under general dynamic state-control-law constraints. We first propose a customized relaxed control formulation to cope with the dynamic mixed constraints and establish the existence of an optimal control using compactification argument in the proper canonical spaces to accommodate singular controls. To further characterize the optimal pair of regular and singular controls, we treat the controlled McKean-Vlasov process as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on canonical spaces with constraints on Banach space, allowing us to derive the stochastic maximum principle (SMP) and a class of constrained BSDE using a new Lagrange multipliers method. Additionally, we investigate the uniqueness and the stability result of the solution to the constrained FBSDE associated with the constrained MFC with singular controls.
Keywords: Mean-field control, singular control, dynamic state-control-law constraint, compactification, stochastic maximum principle, constrained FBSDE
2020 MSC: 60H30; 49N80; 46N10; 60H10
@@ 1 Introduction
To overcome the high complexity in large stochastic system caused by the mutual influence among agents, mean-field game (MFG) problems (see [27] and [31]) have been proposed and popularized in academic studies, which is to find a Nash equilibrium in a competitive game formulation where each agent’s impact on the aggregate distribution is negligible. MFC problems have also gained considerable attention, which concerns the cooperative population where all agents share the same goal to attain the social optimum. The present paper targets at the MFC problem, which is often interpreted as an optimal control of McKean-Vlasov dynamics by a social planner. Some pioneer studies on this topic, just to name a few, can be found in [2, 8, 11, 1, 6] for the SMP method; [30] and [18] for the limit theory; [32] and [35] for the dynamic programming principle (DPP) and HJB equation on the Wasserstein space of probability measures; [4], [36] and [37] for the viscosity solution and comparison principle of the HJB equation. However, it is notable that most existing studies on MFC are limited to regular controls.
@@
The goal of the paper is to contribute to the study of MFC problems by featuring both the singular control and the joint dynamic state-control-law constraints. Motivated by practical risk regulations, liquidity requirements, or feasibility restraints, different types of constraints have been considered in stochastic control problems, but predominantly in single agent’s model. Recently, in mean-field models with only regular controls, there are some emerging research studies to incorporate state and/or law constraints, see [23] for the level-set approach to tackle the law constraint in MFC problems; [15] for the analysis of mean-field limits under state constraints; [16] for the formulation of controlled FP equation under state constraints; [7] for the optimization formulation of the constrained MFC problem and the derivation of the SMP using the generalized Lagrange multipliers method.
@@
The present paper fills the gap in the literature by advancing some theoretical foundations for constrained MFC problems with singular control. The first contribution of the paper is to address the existence of an optimal pair of regular and singular controls by developing the customized relaxed control formulation and martingale problem to embed the dynamic mixed constraints. That is, we embed the dynamic mixed constraints in item (iv) of the admissible control rules in Definition 3.1. We then present in Lemma 3.6 a martingale problem characterization associated to the admissible control rules and establish in Lemma 3.8 the closedness property of the set of probability measures fulfilling the constraints. By developing technical compactification arguments to accommodate the singular control and combining the previous lemmas using the embedded constraints, we are able to show that the set of approximately optimal control rules is compact (see Lemma 3.9), and hence the set of optimal control rules is compact and convex that yields the existence of an optimal relaxed control (see Theorem 3.10). By standard argument, we also obtain the existence of an optimal strict control for the constrained MFC problem with singular control. We emphasize that the classical compactification technique for singular controls in [25] has only been developed in unconstrained settings. In contrast, our work is, to the best of our knowledge, the first to consider compactification under joint constraints over the control, state and law. Rather than directly coping with the dynamic constraints, we transform the problem into constraints over the probability measures on the canonical space. This allows us to extend the relaxed-control framework to the constrained setting. The difficulty caused by constraints is handled by a new result established in Lemma 3.7, which is new to the literature. Moreover, by leveraging the measurable selection result established in Lemma 3.5, we can convert any probability measure on the canonical space into a controlled process that satisfies the original joint dynamic constraints.
@@
Our second contribution is to further characterize the optimal controls by establishing the SMP for the constrained MFC problem with singular control and derive the adjoint constrained BSDE. However, some conventional techniques such as the first order variation or the spike variation of controls may fail in our setting as the mixed constraints might be violated. Inspired by our recent study [7] for constrained MFC with regular controls, we propose to regard the controlled McKean-Vlasov dynamics with singular control as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on some proper canonical spaces with general constraints on Banach space (see the formulation (4)-(19)). Similar to [7], we are allowed to derive the SMP by using the generalized Fritz-John (FJ) optimality condition and the Lagrange multipliers method. To be more precise, when the forward McKean-Vlasov SDE is taken as an equality constraint, the FJ condition for the control variable gives the SMP, and the Lagrange multipliers in the space of Itô processes induced by the FJ condition on the state variable yields the adjoint processes as the solution to the constrained BSDE, see Theorem 4.2. To this end, the core idea is to treat the pair $(X,\alpha)\in\mathbb{M}^{2}\times\mathscr{U}[0,T]$ of the state and control processes as the optimization variable within an abstract constrained optimization problem on Banach spaces. A crucial step is to consider the Hilbert space $\mathbb{M}^{2}$ for the state process $X=(X_{t})_{t\in[0,T]}$. In particular, to cope with singular controls, it is necessary to modify the space $\mathbb{M}^{2}$ by integrating the singular control component (refer to the space $\mathscr{K}$ defined by (4)). Consequently, we must endow the newly defined policy set $\mathscr{K}$ with an appropriate topology, enabling us to utilize the Lagrangian multiplier method proposed in [7] to derive the SMP for our constrained MFC with singular controls.
The third contribution is to analyze the non-standard constrained FBSDE in (5)-(31) associated with our constrained MFC problem with singular control. It is worth noting that the Lagrange multiplier $\eta$ in Theorem 4.2 is treated as part of the solution to the constrained BSDE. The constrained FBSDE (5)-(31) differs substantially from the conventional reflected BSDE that has been extensively studied since [20]. A typical reflected BSDE takes the form given by
$Y_{t}$ $=\xi+\int_{t}^{T}f(s,Y_{s},Z_{s})\mathrm{d}s+K_{T}-K_{t}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s},~~Y_{t}\geq S_{t},~~\int_{0}^{T}(Y_{t}-S_{t})\mathrm{d}K_{t}=0,$
where, $K=(K_{t})_{t\in[0,T]}$ is an adapted increasing process enforcing the constraint. In contrast, in our formulation (Theorem 4.2), the boundary condition for the adjoint process $Y$ is characterized by $\int_{0}^{T}(G(t)^{\top}Y_{t}+c)\mathrm{d}\widehat{\zeta}_{t}=0$, where the singular control $\widehat{\zeta}=(\widehat{\zeta}_{t})_{t\in[0,T]}$ enters the forward dynamics of the state process rather than the backward equation. Moreover, a complementary slackness condition is imposed on the forward component $\mathbb{E}[\int_{0}^{T}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})\eta_{t}^{i}\mathrm{d}t]=0$ for $i\in I$, where the multiplier $\eta^{i}=(\eta_{t}^{i})_{t\in[0,T]}$ appears only in the backward dynamics. This cross-coupled structure leads to a new and more intricate FBSDE system. We then verify the constraint qualification and obtain the KKT optimality condition in Lemma 5.1. As a result, we further establish the uniqueness of the solution to the nonstandard constrained FBSDE in (5)-(31), see Theorem 5.4. Furthermore, under some enhanced assumptions, we also prove the stability result of the solution in the sense of (35) in Theorem 5.5.
@@
The rest of the paper is organized as follows. Section 2 introduces the constrained MFC problem with singular control and standing assumptions imposed on the model. Section 3 gives the relaxed control formulation and establishes the existence of optimal control to the constrained MFC problem by applying compactification arguments. Section 4 formulates the equivalent optimization problem with constraints in Banach space, for which we develop the SMP and adjoint constrained BSDE using Lagrange multipliers method. Lastly, the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control are discussed in Section 5. The proofs of some auxiliary results are provided in Appendix.
@@
${\cal W}_{p,E}(\mu,\nu)$ $p$-order Wasserstein metric between $\mu\in{\cal P}_{p}(E)$ and $\nu\in{\cal P}_{p}(E)$
$\partial_{\mu}f$ Lions derivative of $f:{\cal P}_{2}(\mathbb{R}^{n})\to\mathbb{R}$ with respect to $\mu\in{\cal P}_{2}(\mathbb{R}^{n})$
@@ 2.1 MFC with singular control and mixed constraints
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. For $r,l,k\in\mathbb{N}$ and $p>2$, let $W=(W_{t})_{t\in[0,T]}$ be a standard $r$-dimensional $(\mathbb{P},\mathbb{F})$-Brownian motion and $U\subset\mathbb{R}^{l}$ be a closed and convex (regular) set as the policy space. Define the space of regular controls $\mathscr{U}[0,T]$ as the set of $\mathbb{F}$-progressively measurable processes such that $\mathbb{E}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. For the space of singular controls, define $A^{k}[0,T]$ as the set of functions $a:[0,T]\mapsto\mathbb{R}^{k}$ such that $a=(a_{t}^{i})_{t\in[0,T]}$ is càdlàg and non-decreasing with $a_{0}^{i}\geq 0$ for $i=1,\dots,k$. Let $\mathcal{A}^{k}[0,T]$ be the set of $\mathbb{F}$-progressively measurable processes $\zeta=(\zeta_{t})_{t\in[0,T]}$ with paths in $A^{k}[0,T]$ and $\mathbb{E}[|\zeta_{T}|^{p}]<\infty$.
@@
For the above controlled McKean-Vlasov dynamics, we introduce dynamic state-control-law constraints. Let $d\in\mathbb{N}$ be the number of inequality constraints. For $i\in I:=\{1,\dots,d\}$, consider $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $i\in I$ such that the dynamic state-control-law inequality constraints are described by, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.
@@
$\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}J(\alpha,\zeta)$ (3)
$\qquad:=\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}\mathbb{E}\left[\int_{0}^{T}f(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\mathrm{d}t+{\int_{[0,T]}c(t)d\zeta_{t}+g(X_{T}^{\alpha,\zeta},\mathcal{L}(X_{T}^{\alpha,\zeta})}\right],$
where $f:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$, $c:[0,T]\mapsto\mathbb{R}^{k}$ and $g:\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\to\mathbb{R}$ are respectively the running cost function, the cost function associated with singular controls and the terminal cost function.
@@ 2.2 Standing assumptions
    The mappings $b,\sigma,f,g$ and $\phi^{i}$, for $i\in I$ are Borel measurable, $b,\sigma$ are continuous and $f,g$ is lower semi-continuous (l.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$.
@@
    $-M(1+|x|^{p}+M_{2}(\mu))+M_{1}|u|^{p}$ $\leq f(t,x,\mu,u)\leq M(1+|x|^{p}+M_{2}(\mu)+|u|^{p}),$
    $-M(1+|x|^{p}+M_{2}(\mu))$ $\leq g(x,\mu)\leq-M(1+|x|^{p}+M_{2}(\mu))$
    where $p>2$ is given at the beginning of this section.
@@
    The mappings $b,\sigma,f,g$ and $\phi^{i}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $\phi^{i}$ is jointly continuous in $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$.
@@
    There exists $M>0$ such that $|\delta(t,x,\mu,u)|\leq M(1+|x|+|u|+\sqrt{M_{2}(\mu)})$ for $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ and $\delta\in\{b,\sigma,\nabla_{u}b(\cdot),\nabla_{u}\sigma(\cdot),\nabla_{x}f(\cdot),\nabla_{u}f(\cdot),\nabla_{x}g\}$.
@@
    The mappings $b(t,x,\mu,u),\sigma(t,x,\mu,u),f(t,x,\mu,u),g(x,\mu)$ and $\phi^{i}(t,x,\mu,u)$ are continuously Lions-differentiable w.r.t. $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there exists $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $\left(\int_{\mathbb{R}^{n}}|\partial_{\mu}f(t,x,\mu,u)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}$ $\leq M\left(1+|x|+|u|+\sqrt{M_{2}(\mu)}\right),$
    $\left(\int_{\mathbb{R}^{n}}|\partial_{\mu}g(x,\mu)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}$ $\leq M\left(1+|x|+\sqrt{M_{2}(\mu)}\right)$
@@
    $G:[0,T]\mapsto\mathbb{R}^{n\times k}$ and $c:[0,T]\mapsto[0,\infty)^{k}$ are deterministic continuous functions. Moreover, there exists a constant $c>0$ such that $|G(t)-G(s)|\leq c|t-s|$ and $|G(t)\boldsymbol{\beta}|^{2}\geq c|\boldsymbol{\beta}|$ for any vector $\boldsymbol{\beta}\in\mathbb{R}^{k}$ and $t,s\in[0,T]$.
@@
    The constrained function $\phi$ is independent of the control variable and has the linear form $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t)$ for $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Here, $(A,B)=(A_{t},B_{t})_{t\in[0,T]}\in C([0,T];\mathbb{R}^{n}\times\mathbb{R}^{n})$ and $C=(C(t))_{t\in[0,T]}\in C([0,T];\mathbb{R}^{n})$ and $|A(t)|>0$ and $|A(t)+B(t)|>0$ for $t\in[0,T]$.
@@
    The Hamiltonian $H$ defined by, $H(t,x,\mu,u,y,z):=b(t,x,\mu,u)^{\top}y+\mathrm{tr}(\sigma(t,x,\mu,u)^{\top}z)+f(t,x,\mu,u)$ is strictly convex in $u\in U$. There exists a measurable mapping $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ with at most quadratic growth in $(x,\mu,y,z)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ such that $\hat{u}(t,x,\mu,y,z)\in\mathop{\arg\min}_{u\in U}H(t,x,\mu,u,y,z)$ and $\hat{u}$ is Lipschitz continuous, i.e., there is $M>0$ such that, for $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$,
@@
    Define $F(t,x,\mu,u,y,z):=\nabla_{x}H(t,x,\mu,u,y,z)+\int_{\mathbb{R}^{n}}\partial_{\mu}H(t,x^{\prime},\mu,u,y,z)(x)\mu(\mathrm{d}x^{\prime})$, $\hat{F}(t,x,\mu,y,z):=F(t,x,\mu,\hat{u}(t,x,\mu,y,z),y,z)$ and $\Phi(t,x,\mu,y,z):=(-\hat{F},\hat{b},\hat{\sigma})(t,x,\mu,y,z)$. The following monotone condition holds:
@@
  • (C5)
    Set $\Psi(x,\mu):=\nabla_{x}g(x,\mu)+\int_{\mathbb{R}^{n}}\partial_{\mu}g(x^{\prime},\mu)(x)\mu(\mathrm{d}x^{\prime})$ for $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$ and $\Psi$ satisfies the monotone condition $(\Psi(x_{1},\mu)-\Psi(x_{2},\mu))(x_{1}-x_{2})\geq 0$ and $\int_{\mathbb{R}^{n}}(\Psi(x,\mu_{1})-\Psi(x,\mu_{2}))x^{\prime}(\mu_{1}-\mu_{2})(\mathrm{d}x^{\prime})\geq 0$ for $x,x_{1},x_{2}\in\mathbb{R}^{n}$ and $\mu,\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{n})$.